OpenStudy (solomonzelman):

        ³√(8+h) - 2 Lim ----------- h→0 h I have to identify the f(x) and the "a" (the slope of a line tangent to f(x) at "a", which is in other words a tangent line to f(x) at (a, f(a) )

OpenStudy (solomonzelman):

This is using the \[\large\color{ slate }{\large {\bbox[5pt, yellow ,border:2px solid white ]{ \rm \lim_{h \rightarrow 0} f'(x)=\frac{f(x+h)-f(x)}{h} }}}\]

OpenStudy (anonymous):

(8, 2)

OpenStudy (solomonzelman):

how do I know what f(x) is ?

OpenStudy (anonymous):

\[f \left( x \right)=\sqrt[3]{x}\]

OpenStudy (solomonzelman):

and how does the 8 get there?

OpenStudy (anonymous):

a = 8

OpenStudy (solomonzelman):

How do you do this? Do you like set an equation to solve ? I know how to simplify the limit, but I can't identify the f(x) and the "a" Or is it just one of those common sense questions, because then I guess I am just very stupid.